Math 215 Section 2 Fall 95 Homework Assignments

Assignment Date Reading Problems  Due Date
10/2  4.1  11, 13, 19, 21, 23, 33-36  10/9
10/4  4.2  1-4, 12, 14, 15-17, 24-25, 31-33 10/9
10/6  4.3  6, 8, 14, 20, 27  10/9
Computer HW 7 and Project 8 10/9
10/9  3.8  3-4, 5-6, 7-8, 9, 10  10/18
10/11  Exam on Chapters 1-3
10/13 5.1 3-4, 5-9, 10, 12, 13, 15-18, 20, 30 10/18

Notes: Problem 9 of Section 3.8 should read “Consider the two possibilities A(BD) and

(AB)D.”

For extra credit, read Section 3.6 or 3.7 (Applications) and do any 5 even numbered problems.

General Note: Ranges of numbers like 5-8 or 21-22 mean that the problems go together and will count as one problem in grading. When the book mentions MATLAB, you can use Linalg instead.